# Measuring Social Biases in Masked Language Models by Proxy of Prediction Quality

### v2

@@ 3.4 Recovering Social Biases in Pretrained MLMs
Table 1: MLM predictions and ground truth masked token ranks for input sentences under the ime. The processed input sentences are "The [poor, rich] are really ignorant about how to handle the money they do have", and are contained in $S^{\text{dis}}$ or $S^{\text{adv}}$ from CPS with characteristic words poor and rich respectively. Masked tokens are in bold.
We probe for MLM preferences using the ime, which masks one token at a time until all tokens have been masked, or until we have $n$ logits or predictions for a sentence with $n$ tokens. Appendix I shows an example of the ime for one model and text example.
@@
$\begin{split}\Delta\text{P}(t|s_{\setminus t_{m}};\theta)&=\bigl{(}\log P(t_{p}|s_{\setminus t_{m}};\theta)\\ &\hskip 12.0pt-\log P(t_{m}|s_{\setminus t_{m}};\theta)\bigr{)}\\ &=\Delta\textsc{P}(w;\theta)\end{split}$ (3)
We redefine Equation 1 and propose a modified version $\Delta\text{P}(w)$ as shown in Equation 3.
@@
For a given MLM with parameters $\theta$, we propose attention-weighted measures $\textsc{crra}(w)$ and $\Delta\text{PA}(w)$ defined in Equations 4 and 5 respectively, where $a_{w}$ is the average of all multi-head attentions associated with the ground truth token $w$, and $P(t_{m}|s_{\setminus t_{m}})$ and $\rho(t_{m}|s_{\setminus t_{m}})$ are the probability score and rank of the masked token respectively. We extend these definitions for a sentence $s$ as shown in Equations 6 and 7.
@@
We compute measure $f$, $\forall f\in\{\textsc{crr}_{T}(s),\\ \Delta\textsc{p}_{T}(s),\textsc{crra}_{T}(s),\Delta\textsc{pa}_{T}(s)\}$, where $T$ is an MLM transformer and $s$ is a sentence, $\forall s\in S^{\text{dis}}$ and $\forall s\in S^{\text{adv}}$. [^5] We apply the Shapiro-Wilk test Shapiro and Wilk (1965) for normality to each measure and did not find evidence that the measures were not drawn from a normal distribution. The same was found for the difference of each of these measures between sentence sets relative to the same transformer $T$. Appendix N shows these measures (likelihood scores) for an example sentence $s$ in SS and CPS. [^6] Greater MLM preference based on prediction quality is reflected by crr, crra, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ values closer to 0 (if a sentence with bias against advantaged groups has a greater value relative to its paired counterpart, the MLM is deemed to prefer bias against disadvantaged groups). The opposite is true for measures csps, sss, aul and aula.
We define sets of measures $M_{1}$ and $M_{2}$ ($M_{1}\cap M_{2}=\emptyset$), where $M_{1}=\{\Delta\textsc{p},\textsc{crr},\textsc{crra},\Delta\textsc{pa}\}$ and $M_{2}=\{\textsc{aul},\textsc{aula},\textsc{csps},\textsc{sss}\}$.
$\Delta f_{T}(i)=\begin{cases}f_{T}(S^{\text{adv}}_{i})-f_{T}(S^{\text{dis}}_{i}),&\text{if }f\in M_{1}\\ f_{T}(S^{\text{dis}}_{i})-f_{T}(S^{\text{adv}}_{i}),&\text{if }f\in M_{2}\end{cases}$ (8)
We apply Equation 8 to estimate the preference of a transformer $T$ for $s\in S^{\text{dis}}$ relative to $s\in S^{\text{adv}}$ for paired sentence $s$ and measure $f$, $\forall f\in\{\textsc{crr},\\ \textsc{crra},\Delta\textsc{p},\Delta\textsc{pa},\text{{csps}},\textsc{sss},\text{{aul}},\text{{aula}}\}$. We define a bias score for a pretrained MLM as bspt (Equation 9).
@@
$\mathds{1}$ is a per-model indicator function which returns $1$ if transformer $T$ has a larger preference for $S^{\text{dis}}$ relative to $S^{\text{adv}}$ and 00 otherwise, as estimated for a measure $f$ by Equation 8. bspt represents the proportion of sentences with higher relative bias against disadvantaged groups for a given measure, where values above 50 indicate greater relative bias against disadvantaged groups for an MLM.
3.4.1 Pretrained MLM Bias Scores
We use bspt to compare preferences for a given MLM and report results for all considered measures and MLMs in Appendix K, including corresponding tables and a detailed analysis of results for CPS and SS, focusing on race bias in particular. We report overall bias scores in Table 2.
@@ 3.4.1 Pretrained MLM Bias Scores
Table 2: Overall bias scores for pretrained MLMs using bspt$(f)$ with considered measures $\forall f\in\{\textsc{crr}_{T}(s),\\ \Delta\textsc{p}_{T}(s),\textsc{crra}_{T}(s),\Delta\textsc{pa}_{T}(s)\}$ on CPS and SS, where D- denotes retrained. Color-coded from darkest to lightest, with lower values represented by the darkest shade and higher values by the lightest shade.
@@
We find that only aul and aula estimate overall bias scores below 50, where the corresponding MLM is $\text{BERT}_{\text{unc}}$ in each case.
3.4.2 Alignment with Human Annotated Biases
@@
We compare the alignment (agreement) between measures and bias ratings in CPS. Sentence pairs in CPS received five annotations in addition to the implicit annotation from the writer. We map these sentences to a binary classification task, where a sentence is considered biased if it satisfies criteria from Nangia et al., 2020, where (1) at least three out of six annotators (including the implicit annotation) agree a given pair is socially biased and (2) the majority of annotators who agree a given pair is socially biased agree on the type of social bias being expressed. [^9] This experiment setting gives 58 unbiased pairs and 1,450 biased pairs for binary classification.
@@
We compute evaluation measures derived from MLMs to predict whether a pair is biased or unbiased at varying thresholds. All measures are computed for each sentence in a pair. Thresholds for bias scores computed on sentences with bias against advantaged and disadvantaged groups respectively maximize area under the ROC Curve for that measure. We find that one or more of our measures outperform aul and aula in their agreement with human annotators on CPS for $\text{RoBERTa}_{P}$ and $\text{BERT}_{P,\text{unc}}$ based on higher AUROC values if considered MLMs exhibit bias towards disadvantaged groups (ROC curves shown in Appendix M).
3.5 Recovering Social Biases Introduced by Retraining MLMs
@@
We re-train each transformer under consideration using the PyTorch Python library with P100 and T4 GPUs on cased (RoBERTa and distilRoBERTa) and uncased ($\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$) versions of CPS sentences. [^10] The results are from retraining with 0.15 mlm probability (as in Devlin et al., 2019), using 80 percent of data for training and 20 percent for validation. MLMs were retrained for 30 epochs, reaching minimum validation loss at epoch 30.
@@
Figure 1: Difference between ECDFs for crr distribution for sentences in $S^{\text{dis}}$ and $S^{\text{adv}}$ for re- and pretrained RoBERTa and the race bias category in CPS. The line at $y=0$ separates what is biased against disadvantaged groups on the positive y-axis from what is biased against advantaged groups on the negative y-axis.
@@
Figure 1 shows the difference in ECDFs for the $\textsc{crr}_{m}(s)$ measure on the race bias category in CPS, where $m$ is retrained MLM $\text{RoBERTa}_{R}$. When compared with pretrained RoBERTa (RoBERTa-P), we observe an upwards shift in the difference of ECDFs for the crr difference between sentence sets after retraining RoBERTa on $S^{\text{dis}}$ (RoBERTa-R-DIS), and a downwards shift after retraining on $S^{\text{adv}}$ (RoBERTa-R-ADV). This is expected since retraining MLMs on $S^{\text{dis}}$ or $S^{\text{adv}}$ should shift the MLM preference towards the corresponding bias type. In this sense, the figure illustrates how the difference in ECDF distributions for an MLM can visually represent a contextual shift in relative bias using proposed measure crr.
We define a proxy for the relative preference between two MLMs with the model-comparative indicator function $\mathds{1}$, which returns $1$ if transformer $T_{1}$ has a larger preference for $S^{\text{dis}}$ relative to $S^{\text{adv}}$ than transformer $T_{2}$ and 00 otherwise, as estimated for a measure $f$ by Equation 8. bsrt can be applied to compare pre- and retrained MLMs within the same model class and recover biases introduced by MLM retraining. Values above 50 indicate greater bias against disadvantaged groups for transformer $T_{1}$ relative to $T_{2}$.
3.5.1 Retrained MLM Bias Scores
@@
We compute bsrt for retrained MLMs and all measures and bias categories on CPS and report results in Appendix P. [^11] We assess whether the relative differences in proportions of bias between re- and pretrained transformers are significant according to McNemar’s test McNemar (1947). In general, each measure produces a bias score in accordance with the particular retraining dataset used ($S^{\text{dis}}$ or $S^{\text{adv}}$) for almost all significant results across MLMs, demonstrating that proposed function bsrt can be applied to estimate the bias against disadvantaged groups for transformer $T_{1}$ relative to its pretrained base $T_{2}$.
We find that crr, crra, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ are typically more accurate than csps, aul, and aula for measuring social biases introduced by retraining MLMs, and outperform others with regard to sensitivity for relative changes in MLM bias due to retraining, indicated by larger and smaller scores and more frequently significant relative difference in proportions of bias between re- and pretrained transformers. In every case and across considered MLMs, one or more of our measures reports the highest retraining bias scores for MLMs retrained on $S_{\text{dis}}$ and the lowest for MLMs retrained on $S_{\text{adv}}$.
For $\text{RoBERTa}_{R}$ retrained on $S_{\text{adv}}$ from CPS, each measure gives scores below 50 as expected. However, aul and aula give insignificant results for physical appearance bias.121212crr, crra, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ give significant results below 50 for each bias category with lower scores than other measures in general. For $\text{BERT}_{R}$ retrained on $S_{\text{adv}}$, aul and aula overestimate physical appearance, gender, disability, and socioeconomic biases and give insignificant results for each, while all proposed measures give significant results below 50 for each category as expected. Similarly, for $\text{distilRoBERTa}_{R}$, aula overestimates physical appearance bias and gives insignificant results for physical appearance, gender, and age, aul gives insignificant results for physical appearance, and csps and aul give insignificant results for age.131313aul and aula also give insignificant results for physical appearance, disability, and sexual orientation for $\text{distilBERT}_{R}$ retrained on $S_{\text{adv}}$. Overall, we find that aul, aula, and csps overestimate physical appearance bias introduced by retraining MLMs compared to proposed measures and based on bsrt.
@@
For all MLMs retrained on $S_{\text{dis}}$, crr, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ give results above 50 for each bias category as expected, with higher scores than other measures in almost every case. [^14] This also applies to crra except for disability bias using $\text{distilBERT}_{R}$. In addition, proposed measures are significant for every bias type using $\text{BERT}_{R}$ and $\text{RoBERTa}_{R}$. This is also true for $\text{distilBERT}_{R}$ (besides crra for sexual orientation) and $\text{distilRoBERTa}_{R}$ (besides crra for physical appearance and crr for disability). In contrast, aul, aula, and csps have 5, 6, and 11 insignificant results respectively across all considered MLMs retrained on $S_{\text{dis}}$, and aula and csps each give 3 and 4 bias scores below 50. Notably, aul and crra each give 1 bias score below 50. These are concerning underestimations of biases introduced by retraining MLMs only on sentences with biases against disadvantaged groups from CPS.151515csps, and aula for $\text{RoBERTa}_{R}$ (Appendix P) give scores less than 50 for age bias. aula also gives a score less than 50 for sexual orientation bias. csps, aul, and aula give insignificant results for sexual orientation and age, along with religion and physical appearance for csps and disability for aula.
3.5.2 Alignment with Retraining Biases
We frame a binary classification task where bsrt above 50 indicates increased preference (after retraining) for sentences with bias against disadvantaged groups in CPS (1), and vice versa for scores below 50 (0). [^16] There are 72 predictions per measure across MLMs and bias types. Half of binary truths are 1 and 0 respectively since MLMs retrained on $S_{\text{dis}}$ should score above 50 for all bias types (vice versa for MLMs retrained on $S_{\text{adv}}$) and $S_{\text{dis}}$ has the same length (number of sentences) as $S_{\text{adv}}$.
Re-train Dataset: $\forall s\in s^{dis}$ for CPS

### v3

@@ 3.1 Biases in Pretrained MLMs
Table 1: MLM predictions and ground truth masked token ranks for input sentences under the ime. The processed input sentences are "The [poor, rich] are really ignorant about how to handle the money they do have", and are contained in $S_{\text{dis}}$ or $S_{\text{adv}}$ from CPS with characteristic words poor and rich respectively. Masked tokens are in bold.
We probe for MLM preferences using the ime, which masks one token at a time until all tokens have been masked, or until we have $n$ logits or predictions for a sentence with $n$ tokens. Appendix I shows an ime example for one model and text.
@@
$\begin{split}\Delta\text{P}(t|s_{\setminus t_{m}};\theta)&=\bigl{(}\log P(t_{p}|s_{\setminus t_{m}};\theta)-\log P(t_{m}|s_{\setminus t_{m}};\theta)\bigr{)}\\ &=\Delta\textsc{P}(w;\theta)\end{split}$ (3)
Motivated by the typically long-tail distribution of probabilities output by MLMs, we redefine Equation 1 and propose a modified version $\Delta\text{P}(w)$ as shown in Equation 3. The log transformation reduces sensitivity to differences when probabilities are high but enhances relative differences when probabilities are low, improving sensitivity to subtle but potentially meaningful variations in low-probability token choices.
@@
For a given MLM with parameters $\theta$, we propose attention-weighted measures $\textsc{crra}(w)$ and $\Delta\text{PA}(w)$ defined in Equations 4 and 5 respectively, where $a_{w}$ is the average of all multi-head attentions associated with the ground truth token $w$ [^6] Averaging the multi-head attentions associated with a ground truth token provides a potentially balanced representation of token-level contributions by capturing diverse contextual relationships encoded by individual heads while reducing noise from any single head., and $P(t_{m}|s_{\setminus t_{m}})$ and $\rho(t_{m}|s_{\setminus t_{m}})$ are the probability score and rank of the masked token respectively. We extend these definitions for a sentence $s$ as shown in Equations 6 and 7.
@@
[…]
$\Delta f_{T}(i)=\begin{cases}f_{T}(S_{\text{adv}}(i))-f_{T}(S_{\text{dis}}(i)),&\text{if }f\in M_{1}\\ f_{T}(S_{\text{dis}}(i))-f_{T}(S_{\text{adv}}(i)),&\text{if }f\in M_{2}\end{cases}$ (8)
[…]
@@
$\mathds{1}$ is a per-model indicator function which returns $1$ if transformer $T$ has a larger preference for $S_{\text{dis}}$ relative to $S_{\text{adv}}$ and 00 otherwise, as estimated for a measure $f$ by Equation 8. bspt represents the proportion of sentences with higher relative bias against disadvantaged groups for a given measure, where values above 50 indicate greater relative bias against disadvantaged groups for an MLM.
3.2 Biases Introduced by MLM Retraining
We define a proxy for the relative preference between two MLMs with model-comparative indicator function $\mathds{1}$, which returns $1$ if transformer $T_{1}$ has a larger preference for $S_{\text{dis}}$ relative to $S_{\text{adv}}$ than transformer $T_{2}$ and 00 otherwise, as estimated for a measure $f$ by Equation 8. bsrt can be applied to compare pre- and retrained MLMs within the same model class and recover biases introduced by MLM retraining. Values above 50 indicate greater bias against disadvantaged groups for transformer $T_{1}$ relative to $T_{2}$.
4 Experiments and Findings
4.1 Benchmark Datasets for Social Bias
The Crowdsourced Stereotype Pairs Benchmark (CPS; Nangia et al., 2020) contains biased sentences towards historically advantaged and disadvantaged groups along nine forms of social biases. StereoSet (SS; Nadeem et al., 2021) contains intrasentence and intersentence (with context) pairs for four forms of social biases, using the likelihood of modified tokens given unmodified token contexts as proxy for MLM preference. Similarly, CPS contains characteristic words that distinguish sentences within pairs and define the nature of a particular bias towards either advantaged or disadvantaged groups, but instead uses the relative likelihood of unmodified tokens being chosen by the MLM given a modified context (characteristic word) across a sentence pair (see Appendix A for details on bias categories and sentence counts).
To probe for biases of interest that are encoded in MLMs, the scope of our experiments include all bias categories and sentences pairs in CPS and intrasentence pairs in SS, since intersentence pairs are not masked for bias evaluation (Kaneko and Bollegala, 2022). We estimate MLM preference towards a stereotypical sentence over a less stereotypical one for each bias category in CPS and SS.
4.2 Retraining Dataset
CPS provides a more diverse alternative to biases expressed by sentence pairs in SS. Biases widely acknowledged in the United States are well represented in CPS [^9] CPS categories are a ”narrowed” version of the US Equal Employment Opportunities Commission’s list of protected categories Nangia et al. (2020). compared to SS, and there is greater diversity of sentence structures in CPS Nangia et al. (2020). CPS has been found to be a more reliable benchmark for pretrained MLM bias measurement than SS, and the validation rate of CPS is 18% higher than SS. Nangia et al. (2020). Based on these findings, paired with (1) the computational expense and time-consumption involved with retraining MLMs under the MLMO and (2) concerns regarding standard masked language modeling metric viability on SS, we proceed to use sentence sets in CPS to re-train MLMs and validate methods for estimating the biases that are introduced Nangia et al. (2020) (see Appendix A.2 for details). We re-train MLMs $\forall s\in S_{\text{dis}}$ or $\forall s\in S_{\text{adv}}$, where $s$ is a sentence in CPS biased towards either advantaged ($S_{\text{adv}}$) or disadvantaged groups ($S_{\text{dis}}$).
4.3 Transformer-based Language Models
We denote pre- and retrained transformers as $T_{P}$ and $T_{R}$ respectively. The subscript ${}_{\text{unc}}$ denotes an uncased model. We report results from the following transformer-based language models available through the HuggingFace library Wolf et al. (2020): bert-base-uncased ($\text{BERT}_{\text{unc}}$; Devlin et al., 2019), roberta-base (RoBERTa; Liu et al., 2019b), distilbert-base-uncased ($\text{distilBERT}_{\text{unc}}$; Sanh et al., 2019), and distilroberta-base (distilRoBERTa; Liu et al., 2019a).
4.4 Pretrained MLM Bias Scores
We use bspt to compare MLM preferences, report overall bias scores in Table 2, and provide detailed results for all measures and MLMs in Appendix K.
@@ 4.4 Pretrained MLM Bias Scores
[…]
@@
4.4.1 Human Annotated Bias Alignment
@@
We compare the alignment (agreement) between measures and bias ratings in CPS. Sentence pairs in CPS received five annotations in addition to the implicit annotation from the writer. We map sentences to a binary classification task, where a sentence is considered biased if it satisfies criteria from Nangia et al., 2020, where (1) at least three out of six annotators (including the implicit annotation) agree a given pair is socially biased and (2) the majority of annotators who agree a given pair is socially biased agree on the type of social bias being expressed. [^11] This experiment setting gives 58 unbiased pairs and 1,450 biased pairs for binary classification.
@@
We compute evaluation measures derived from MLMs to predict whether a pair is biased or unbiased at varying thresholds. All measures are computed for each sentence in a pair. Thresholds for bias scores computed on sentences with bias against advantaged and disadvantaged groups respectively maximize area under the ROC Curve for that measure. We find one or more of proposed measures $\Delta\textsc{pa}$ and crra, along with modified $\Delta\textsc{p}$ and crr, outperform aul and aula in their agreement with human annotators on CPS for $\text{RoBERTa}_{P}$ and $\text{BERT}_{P,\text{unc}}$ based on higher AUROC values if MLMs exhibit bias towards disadvantaged groups (ROC curves in Appendix M).
4.5 Retrained MLM Bias Scores
Model Retrain dataset $f:\min/\max\textsc{bsrt}\ \exists c\in C$ $f:p<a\ \forall c\in C$
$\text{RoBERTa}_{R}$ $\forall s\in S_{\text{dis}}$ $\max{\textsc{bsrt}}:\{\textbf{$\Delta\textsc{p}$},\textbf{$\Delta\textsc{pa}$}\}$ crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$
$\forall s\in S_{\text{adv}}$ $\min{\textsc{bsrt}}:\{\textbf{{crra}},\textbf{$\Delta\textsc{p}$}\}$ csps, crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$
$\text{distilRoBERTa}_{R}$ $\forall s\in S_{\text{dis}}$ $\max{\textsc{bsrt}}:\{\textbf{{crr}},\textbf{$\Delta\textsc{p}$},\textbf{$\Delta\textsc{pa}$}\}$ $\Delta\textsc{p}$, $\Delta\textsc{pa}$
$\forall s\in S_{\text{adv}}$ $\min{\textsc{bsrt}}:\{\textbf{{crra}},\textbf{$\Delta\textsc{p}$},\textbf{$\Delta\textsc{pa}$}\}$ crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$
$\text{BERT}_{R}$ $\forall s\in S_{\text{dis}}$ $\max{\textsc{bsrt}}:\{\textbf{$\Delta\textsc{p}$},\textbf{$\Delta\textsc{pa}$}\}$ crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$
$\forall s\in S_{\text{adv}}$ $\min{\textsc{bsrt}}:\{\textbf{{crra}},\textbf{$\Delta\textsc{p}$},\textbf{$\Delta\textsc{pa}$}\}$ csps, crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$
$\text{distilBERT}_{R}$ $\forall s\in S_{\text{dis}}$ $\max{\textsc{bsrt}}:\{\textbf{{aul}},\textbf{{crr}},\textbf{$\Delta\textsc{p}$},\textbf{$\Delta\textsc{pa}$}\}$ aul, aula, $\Delta\textsc{p}$, $\Delta\textsc{pa}$
$\forall s\in S_{\text{adv}}$ $\min{\textsc{bsrt}}:\{\textbf{{crra}},\textbf{$\Delta\textsc{p}$},\textbf{$\Delta\textsc{pa}$}\}$ csps, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$
Table 3: Measures $\forall f\in\{\textsc{crr},\textsc{crra},\Delta\textsc{p},\Delta\textsc{pa},\textsc{csps},\textsc{aul},\textsc{aula}\}$ with at least one $\min\textsc{bsrt}$ or $\max\textsc{bsrt}$ across bias categories in CPS for MLMs retrained on $S_{\text{adv}}$ or $S_{\text{dis}}$ respectively (bsrt; Equation 10). $f:p<a\ \forall c\in C$ indicates that, for a measure $f$ and every bias category in CPS, the relative difference in proportions of bias between pre- and retrained transformers is statistically significant according to McNemar’s test (p-value < 0.05).
@@ 4.5 Retrained MLM Bias Scores
We re-train each transformer under consideration using the PyTorch Python library with P100 and T4 GPUs on cased (RoBERTa and distilRoBERTa) and uncased ($\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$) versions of CPS sentences (see Appendix P.1 for more details).
@@
Figure 1: Difference between ECDFs for crr distribution for sentences in $S_{\text{dis}}$ and $S_{\text{adv}}$ for pre- and retrained RoBERTa and the race bias category in CPS. The line at $y=0$ separates what is biased against disadvantaged groups on the positive y-axis from what is biased against advantaged groups on the negative y-axis.
@@
Figure 1 shows the difference in ECDFs for the $\textsc{crr}_{m}(s)$ measure on the race bias category in CPS, where $m$ is retrained MLM $\text{RoBERTa}_{R}$. It illustrates how the difference in ECDF distributions for an MLM can visually represent a contextual shift in relative bias. When compared with pretrained RoBERTa (RoBERTa-P), we observe an upwards shift in the difference of ECDFs for the crr difference between sentence sets after retraining RoBERTa on $S_{\text{dis}}$ (RoBERTa-R-DIS), and a downwards shift after retraining on $S_{\text{adv}}$ (RoBERTa-R-ADV). This is expected since retraining on $S_{\text{dis}}$ or $S_{\text{adv}}$ should shift the MLM preference towards the corresponding bias type.
@@
We compute bsrt for retrained MLMs and all measures and bias categories on CPS and report detailed results in Appendix P.2.121212We assess whether the relative differences in proportions of bias between pre- and retrained transformers are statistically significant according to McNemar’s test McNemar (1947) (p-value ¡ 0.05), using binarized outcomes for bias as given by $f_{R}(S_{\text{adv}})>f_{R}(S_{\text{dis}})$ and $f_{P}(S_{\text{adv}})>f_{P}(S_{\text{dis}})$ to create contingency tables of outcome pairings between pre- and retrained transformers to test for marginal homogeneity. In general, each measure produces a bias score in accordance with the particular retraining dataset used for almost all significant results across MLMs, demonstrating that proposed function bsrt can be applied to estimate the bias against disadvantaged groups for transformer $T_{1}$ relative to its pretrained base $T_{2}$.
In Table 3, we report (1) measures with at least one $\min\textsc{bsrt}$ or $\max\textsc{bsrt}$ across bias categories in CPS for MLMs retrained on $S_{\text{adv}}$ or $S_{\text{dis}}$ respectively, and (2) measures for which the relative difference in proportions of bias between pre- and retrained transformers is statistically significant according to McNemar’s test (p-value < 0.05) for every bias category in CPS.
@@
For all MLMs retrained on $S_{\text{dis}}$, crr, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ give results above 50 for each bias category as expected, with higher scores than other measures in almost every case. crra also gives results above 50 for each bias category except for disability bias using $\text{distilBERT}_{R}$. Similarly, for all MLMs retrained on $S_{\text{adv}}$, crr, crra, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ give results below 50 for each bias category as expected, with lower scores than other measures in almost every case. In contrast, aula and csps give 3 and 4 bias scores below 50 respectively across all MLMs retrained on $S_{\text{dis}}$. Notably, aul and crra give 1 bias score below 50. aul and aula give 4 and 5 bias scores above 50 respectively across all MLMs retrained on $S_{\text{adv}}$.
As shown in Table 3, in every case and across MLMs, one or more of proposed measures reports the highest retraining bias scores for MLMs retrained on $S_{\text{dis}}$ and the lowest for MLMs retrained on $S_{\text{adv}}$. In addition, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ give significant results for every bias type across all MLMs retrained on $S_{\text{dis}}$ or $S_{\text{adv}}$, and crr and crra give significant results for every bias type using $\text{BERT}_{R}$ and $\text{RoBERTa}_{R}$. In contrast, aul, aula, and csps have 5, 6, and 11 insignificant results respectively across all MLMs retrained on $S_{\text{dis}}$, and 1, 12, and 13 insignificant results respectively across all MLMs retrained on $S_{\text{adv}}$.
We find $\Delta\textsc{pa}$, $\Delta\textsc{p}$, and crra show greater sensitivity than crr, csps, aul, and aula for relative changes in MLM bias due to retraining, indicated by larger and smaller scores and more frequently significant relative difference in proportions of bias between pre- and retrained transformers. We find csps, aul, and aula produce concerning underestimations of biases introduced by retraining MLMs only on sentences with biases against disadvantaged groups from CPS, and overestimations from retraining on sentences with biases against advantaged groups (see Appendix P.2 for more details).
4.5.1 Retraining Bias Alignment
Retrain dataset: $\forall s\in S^{dis}$